# Electron-phonon coupling in magnetic materials using the local spin density approximation

### v1

@@ VI Conclusions
We explore the role of electron-phonon interactions in typical magnetic metals. Namely ferromagnetic Fe and Ni where we show that neither the deformation potential nor the electron-phonon coupling present the same strength for the two spin channels. The latter being dominated by Fermi surface nesting effects and the partial density of states of each spin channel. Additionally, we show that the electrical resistivity of Fe and Ni presents important contributions from electron-phonon scattering at room temperature although the relative importance is material dependent. While in Fe the resistivity due to electron-phonon interactions is the dominant contribution at room temperature, for Ni it is less than one-third of the total. Additionally, we show that including the spin degree of freedom to compute resistivity is crucial. Neglecting it can lead to unstable phonons such as in the case of Fe, can overestimate resistivity, and can yield a completely erroneous interpretation for the source of the resistivity. Furthermore, we explore the phonon-mediated superconducting mechanism in these two elemental compounds, a subject rarely addressed with proper numerical calculations. We find suppressed superconductivity, showing that even in a non-magnetic configuration their critical temperatures would be extremely low. These results validate and extend the capabilities of EPW to compute electron-phonon quantities with collinear magnetism. Beyond the methodological advances, our findings highlight how magnetism critically shapes resistive losses and superconductivity in metals that are central to energy applications. By enabling accurate modeling of transport in magnetic systems, this work provides a foundation for the design of low-energy-consumption spintronic devices, improved conductive pathways with reduced losses, and future magnetic materials for energy conversion and transmission. Future extensions to include spin–orbit coupling and non-collinear magnetism will further expand the reach of this framework, opening new directions for first-principles studies of complex magnetic materials.

### v2

@@ VI Conclusions
We explore the role of electron-phonon interactions in typical magnetic metals. Namely, ferromagnetic Fe and Ni, where we show that neither the deformation potential nor the electron-phonon coupling constant present the same strength for the two spin channels, with the latter being dominated by Fermi surface nesting effects and the partial density of states of each spin channel. Additionally, we show that the electrical resistivity of Fe and Ni exhibits important contributions from electron-phonon scattering at room temperature although the relative importance is material dependent. While in Fe the resistivity due to electron-phonon interactions is the dominant contribution at room temperature, for Ni it is less than one-third of the total. Additionally, we show that including the spin degree of freedom to compute resistivity is crucial. Neglecting it can lead to unstable phonons such as in the case of Fe, can overestimate resistivity, and can yield a completely erroneous interpretation for the source of the resistivity. Furthermore, we explore the phonon-mediated superconducting mechanism in these two elemental compounds, a subject rarely addressed with proper numerical calculations. We find suppressed superconductivity, showing that even in a non-magnetic configuration their critical temperatures would be extremely low. These results validate and extend the capabilities of EPW to compute electron-phonon quantities with collinear magnetism. Beyond the methodological advances, our findings highlight how magnetism critically shapes resistive losses and superconductivity in metals that are central to energy applications. By enabling accurate modeling of transport in magnetic systems, this work provides a foundation for the design of low-energy-consumption spintronic devices, improved conductive pathways with reduced losses, and future magnetic materials for energy conversion and transmission. Future extensions to include spin–orbit coupling and non-collinear magnetism will further expand the reach of this framework, opening new directions for first-principles studies of complex magnetic materials.